Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 8:13 Wednesday 23 November 2022 (57864) Print this report
Operator Report - Initial status

ITF State: DOWN
Quick Summary: IMC and RFC locked; all the suspension loops closed with the exception of WI ( WI test mass ); SNEB, SQB2 loops open; Timing error for WI DSP (experts informed since long time); WI large valve and WI-BS valve closed.
Activities ongoing since this morning: none.

Images attached to this report
Comments to this report:
Boschi - 18:03 Wednesday 23 November 2022 (57878) Print this report

Timing errors on both Sa_WI and Sc_WI have been fixed. A minor error on WI TANGO servers was also fixed: the database was not updated with the PSD board replaced in March (current serial #44192).

Search Help
×

Warning

×